# Landau.bib

### v1

@@ 3.3 Variational characterisation for the fuzzy Landau equation
$\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}|{\min\{f\log f,0\}}|{\,\rm d}x{\,\rm d}v$
[…]
where we choose $\kappa>0$ small such that $(\langle x\rangle^{2}+\langle v\rangle^{2})^{-\frac{1-\kappa}{\kappa}}\in L^{1}(\mathbb{R}^{3}\times\mathbb{R}^{3})$. Hence, we have
@@
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{12}}\tilde{\nabla}\log(f_{t})U_{t}{\,\rm d}\eta{\,\rm d}t\geq-\frac{1}{2}\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t-\frac{1}{2}\int_{0}^{T}\mathcal{A}(f_{t},U_{t}){\,\rm d}t.$
@@
$\mathcal{H}(f_{T})-\mathcal{H}(f_{0})=\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{12}}\tilde{\nabla}\log(f_{t})U_{t}{\,\rm d}\eta{\,\rm d}t$
$\geq-\frac{1}{4}\int_{0}^{T}\int_{\mathbb{R}^{12}}\big{(}-\tilde{\nabla}\log(f_{t})\big{)}^{2}ff_{*}\kappa{\,\rm d}\eta{\,\rm d}t-\frac{1}{4}\int_{0}^{T}\int_{\mathbb{R}^{12}}\frac{|U_{t}|^{2}}{{ff_{*}\kappa}}{\,\rm d}\eta{\,\rm d}t.$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{12}}\big{|}\tilde{\nabla}\log(f_{t})\sqrt{ff_{*}\kappa}+\frac{U_{t}}{\sqrt{ff_{*}\kappa}}\big{|}^{2}{\,\rm d}\eta{\,\rm d}t=0.$
@@
$U=-ff_{*}\kappa\tilde{\nabla}\log f$
almost everywhere on $[0,T]\times\mathbb{R}^{12}$. The weak formulation (2.14) for the fuzzy Landau equation holds since $(f,-\kappa ff_{*}\tilde{\nabla}\log f)\in\mathcal{TGRE}_{T}$.
If $f_{t}$ is a $\mathcal{H}$-solutions of (1.1), (notice that the entropy dissipation $\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t<+\infty$ by definition), we take $U=-\kappa ff_{*}\tilde{\nabla}\log f$, then $(f_{t},U_{t})\in\mathcal{TGRE}_{T}$ and
@@
4 Variational characterisation on $\mathbb{R}^{2d}$
Let $d\geq 2$. In this section, in Theorem 4.1, we establish the GENERIC variational characterisation for the fuzzy Landau equation (1.1) on $\mathbb{R}^{2d}$ for a class of kernels $\bar{A}$ and $\kappa$ that satisfy Assumption 4.1 below. These kernels have also been studied for homogeneous/fuzzy Boltzmann equations [erbar2023gradient, EH24].
Assumption 4.1 (kernel assumptions).
The interaction kernel $\bar{A}:\mathbb{R}^{d}\to\mathbb{R}_{+}$ is measurable, continuous, and symmetric
$\bar{A}(v-v_{*})=\bar{A}(v_{*}-v)\quad\forall v,\,v_{*}\in\mathbb{R}^{d}.$
Moreover, for $\gamma\in(-\infty,3]$, the following bounds hold
$C_{\bar{A}}^{-1}\langle v-v_{*}\rangle^{\gamma}\leq\bar{A}(v-v_{*})\leq C_{\bar{A}}\langle v-v_{*}\rangle^{\gamma},$
for some constant $C_{\bar{A}}>0$.
The spatial kernel $\kappa$ satisfies the assumption (2.2)
$\kappa(x-x_{*})={k_{1}}\exp(-{k_{2}}\langle x-x_{*}\rangle),$
Compared to the kernel $A(v-v_{*})=|v-v_{*}|^{2+\gamma}$ in Assumption 2.1, the key point is that the following bounds hold uniformly for a sequence of mollifiers for all $\beta\in(0,1)$
$\bar{A}*_{v}M_{\beta}\leq C\bar{A}$ $\text{and}\quad\bar{A}^{-1}*_{v}M_{\beta}\leq C\bar{A}^{-1},$ (4.1)
where $C>0$ is a constant independent of $\beta$. The proof is analogue to (2.9).
We define the Landau gradient
~$\tilde{\nabla}_{L}f(x,v)=\Pi_{(v-v_{*})^{\perp}}\big{(}\nabla_{v}f(x,v)-\nabla_{v_{*}}f(x_{*},v_{*})\big{)}.$
Notice that the gradient ~$\tilde{\nabla}_{L}$ is not weighted by the kernel $\bar{A}$ compared to $\tilde{\nabla}$ defined as in (1.11)
~$\tilde{\nabla}f(x,v)=\sqrt{A}\Pi_{(v-v_{*})^{\perp}}\big{(}\nabla_{v}f(x,v)-\nabla_{v_{*}}f(x_{*},v_{*})\big{)}.$
According to the integration by parts formula (1.12), we obtain the following expression for the divergence operator associated to ~$\tilde{\nabla}_{L}$:
~$\tilde{\nabla}_{L}\cdot B=\nabla_{v}\cdot\Big{(}\int_{\mathbb{R}^{2d}}\Pi_{(v-v_{*})^{\perp}}\big{(}B(x,x_{*},v,v_{*})-B(x_{*},x,v_{*},v)\big{)}{\,\rm d}x_{*}{\,\rm d}v_{*}\Big{)}.$
Then the entropy dissipation (1.10) can be accordingly written as
$\mathcal{D}_{L}(f)=\frac{1}{2}\int_{\mathbb{R}^{4d}}\bar{A}\kappa ff_{*}|\tilde{\nabla}_{L}\log f|^{2}{\,\rm d}\eta_{d},$
where ${\,\rm d}\eta_{d}$ denotes the Lebesgue measure on $\mathbb{R}^{4d}$.
The GENERIC building block $\{\mathsf{L},\,\mathsf{E},\,\mathsf{M},\,\mathsf{S}\}$ in Appendix A is still a building block if we replace $\mathsf{M}$ by
$\mathsf{M}(f)g=-\frac{1}{2}\tilde{\nabla}_{L}\cdot\Big{(}{\bar{A}}\kappa ff_{*}\tilde{\nabla}_{L}g\Big{)}.$
We define the transport grazing rate equation
$\partial_{t}f+v\cdot\nabla_{x}f+\frac{1}{2}\tilde{\nabla}_{L}\cdot U=0,$ (4.2)
and the curve action
$\mathcal{A}_{L}(f,U)=\frac{1}{2}\int_{\mathbb{R}^{4d}}\frac{|U|^{2}}{ff_{*}\kappa{\bar{A}}}{\,\rm d}\eta_{d}.$
Parallel to Theorem 1.1, we have the following variational characterisation theorem.
Theorem 4.1.
Let the kernels $({\bar{A}},\kappa)$ satisfy Assumption 4.1. Let $(f_{t},\mathcal{U}_{t})$ be solutions to (4.2). Let $(f_{t})_{t\in[0,T]}\subset\mathcal{P}_{2,2+\gamma_{+}}(\mathbb{R}^{2d})$ be a curve satisfying $\int_{0}^{T}\mathcal{E}_{2,2+\gamma_{+}}(f_{t}){\,\rm d}t<+\infty$ and initial entropy $|\mathcal{H}(f_{0})|<+\infty$. Then we have
$\mathcal{J}_{T}(f,\mathcal{U}):=\mathcal{H}(f_{T})-\mathcal{H}(f_{0})+\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t+\int_{0}^{T}\mathcal{A}_{L}(f_{t},\mathcal{U}_{t}){\,\rm d}t\geq 0.$
Moreover, we have $\mathcal{J}_{T}(f,\mathcal{U})=0$ if and only if $f$ is a $\mathcal{H}$-solution for the fuzzy Landau equation (1.1) with the kernels $(\bar{A},\kappa)$.
The proof follows Section 3 analogy. The key is to show the chain rule (3.10) in Theorem 3.6
$\frac{d}{dt}\mathcal{H}(\mu_{t})=\frac{1}{2}\int_{\mathbb{R}^{4d}}\tilde{\nabla}_{L}\log f_{t}{\,\rm d}\mathcal{U}_{t},$
where $\mu_{t}=f_{t}\mathcal{L}$ satisfies the same assumptions as in Theorem 4.1.
We follow the three-step proof for Theorem 3.6. One can repeat the four-step regularisation in Section 3.2.1. The main difference is the collision term in the approximated chain rule (3.32) in Section 3.2.2
$\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{t})-\mathcal{H}(f^{\alpha,\beta,\delta,\theta}_{s})$ (4.3)
$=-\int_{s}^{t}\int_{\mathbb{R}^{2d}}\mathsf{E}^{\alpha,\beta,\delta,\theta}_{r}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}x{\,\rm d}v{\,\rm d}r$
$\quad+\frac{1}{2}\int_{s}^{t}\int_{\mathbb{R}^{4d}}U^{\alpha,\beta,\delta,\theta}_{r}\cdot\tilde{\nabla}\log f^{\alpha,\beta,\delta,\theta}_{r}{\,\rm d}\eta_{d}{\,\rm d}r,$
where the key point is the definition of the unweighted gradient ~$\tilde{\nabla}_{L}$ implying the following identity
(~$\big{(}\tilde{\nabla}_{L}\cdot\mathcal{U}^{\alpha,\delta,\theta}\big{)}*_{v,v_{*}}M_{\beta}=\tilde{\nabla}_{L}\cdot\mathcal{U}^{\alpha,\beta,\delta,\theta}.$
So in Section 3.2.3, for passing to the limit, we only need to treat $\beta\to 0$ limit for the collision term in (4.3) differently. For simplification, we only keep $\beta$ parameter. By Cauchy–Schwarz inequality, we have
$\big{|}U^{\beta}\cdot\tilde{\nabla}_{L}\log f^{\beta}\big{|}\leq\frac{1}{2}\frac{|U^{\beta}|^{2}}{f^{\beta}f^{\beta}_{*}\kappa{\bar{A}}}+\frac{1}{2}|\tilde{\nabla}_{L}\log f^{\beta}|^{2}f^{\beta}f^{\beta}_{*}\kappa{\bar{A}}.$
The limits of $\frac{|U^{\beta}|^{2}}{f^{\beta}f^{\beta}_{*}\kappa{\bar{A}}}$ and |~$|\tilde{\nabla}_{L}\log f^{\beta}|^{2}f^{\beta}f^{\beta}_{*}\kappa{\bar{A}}$ follow analogue to Lemma 3.10. In particularly, by using of the convexity of $(a,b)\mapsto\frac{|a|^{2}}{b}$, we have
${\bar{A}}\kappa\frac{\big{|}\Pi_{(v-v_{*})^{\perp}}\bar{\nabla}(f^{\beta}f^{\beta}_{*})\big{|}^{2}}{f^{\beta}f^{\beta}_{*}}\leq{\bar{A}}\kappa\Big{(}\frac{\big{|}\Pi_{(v-v_{*})^{\perp}}\bar{\nabla}(ff_{*})\big{|}^{2}}{ff_{*}}\Big{)}*_{v,v_{*}}M_{\beta}.$
Then the uniform bounds of $\bar{A}$ in (4.1) ensures that
$\int_{0}^{T}\mathcal{D}(f^{\alpha,\beta}){\,\rm d}t\leq C\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t.$
Then by using the dominated convergence Theorem 3.12, one can pass to the limit by letting $\beta\to 0$ in (4.3).
The case of kernels satisfying Assumption 4.1 can be seen as the cases in Assumption 2.1 with $2+\gamma\geq 0$. Compared to the curve Assumption 2.2, we do not need $L^{p}$ integrability in $v$ since we do not need to control the Fisher information as in Lemma 2.6. On the other hand, we assume a higher moment bound of order $2+\gamma_{+}$ compared to the kernel of order $\gamma_{+}$. The higher moments are used to control the integrability of the grazing rate function $U$ (compared to Lemma 3.5)
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}|U|\langle v\rangle{\,\rm d}\eta_{d}{\,\rm d}t$
[…]
5 Conclusion and Outlook
@@
In this paper, we have rigorously established the GENERIC variational characterisation for the fuzzy Landau equation (1.1) for variants of interaction and spatial kernels. Our results relies on delicate regularisation and error estimates analysis to verify the chain rule for the grazing transport equation. This paper opens up a couple of interesting avenue for future research. First, together with previous works [carrillo2024landau, erbar2023gradient, EH24] on the Boltzmann/Landau equations, our results cast the spatial homogeneous, fuzzy (delocalised), and classical (at least formally) versions of these equations into a common mathematical framework. This paves the way for deriving one equation from another by passing to the limit a certain parameter (such as from the Boltzmann equations to the Landau equations in the grazing limit, or from the fuzzy models to the classical ones in the regularisation limits) by generalising the well-established evolutionary $\Gamma$-convergence methods for gradient flows [sandier2004gamma, mielke2016evolutionary] to GENERIC systems. The second interesting question is to understand relations between the GENERIC variational solutions with existing concepts of solutions such as $\mathcal{H}$-solutions and weak-solutions, and to construct time discrete, iterative variational schemes (for instance, the Jordan- Kinderlehrer-Otto [JKO98] scheme) to numerically solve these equations. Another natural problem is to generalise this framework to other (delocalised) kinetic models such as those discussed in [FZF24]. We will return to these topics in future publications.
@@ Appendix A GENERIC formulation
We consider the space $\mathsf{Z}$ to be the space that consists of all Schwartz functions $f\in\mathcal{S}(\mathbb{R}^{3}\times\mathbb{R}^{3})$ such that $f\mathcal{L}\in\mathcal{P}(\mathbb{R}^{3}\times\mathbb{R}^{3})$. The space is endowed with the $L^{2}$-inner product $\langle f,g\rangle=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}fg{\,\rm d}x{\,\rm d}v$. We formally consider the tangent and co-tangent space at $f\in\mathsf{Z}$ as $T_{f}\mathsf{Z}=\{\xi\in L^{2}(\mathbb{R}^{3}\times\mathbb{R}^{3})\mid\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\xi{\,\rm d}x{\,\rm d}v=0\}$ and $T_{f}^{*}\mathsf{Z}=L^{2}(\mathbb{R}^{3}\times\mathbb{R}^{3})\}$.
@@
$\mathsf{E}(f)=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}\frac{|v|^{2}}{2}f{\,\rm d}x{\,\rm d}v\quad\text{and}\quad\mathsf{S}(f)=-\mathcal{H}(f).$
@@
$\mathsf{M}(f)g$ $=-\frac{1}{2}\tilde{\nabla}\cdot\Big{(}\kappa ff_{*}\tilde{\nabla}g\Big{)}\quad\text{and}$
$\mathsf{L}(f)g$ $=-\nabla\cdot(f\mathsf{J}\nabla g),\quad\mathsf{J}=\begin{pmatrix}0&\mathsf{id}_{3}\\ -\mathsf{id}_{3}&0\end{pmatrix}$
@@
$\langle\mathsf{L}(f)g,g\rangle=\int_{\mathbb{R}^{3}\times\mathbb{R}^{3}}f\nabla g\mathsf{J}\nabla g=0$
@@
$\langle\mathsf{M}(f)g,h\rangle$ $=-\frac{1}{2}\int_{\mathbb{R}^{12}}\tilde{\nabla}\cdot\Big{(}\kappa ff_{*}\tilde{\nabla}g\Big{)}h{\,\rm d}\eta$
$=\frac{1}{2}\int_{\mathbb{R}^{12}}\kappa ff_{*}\tilde{\nabla}g\cdot\tilde{\nabla}h{\,\rm d}\eta=\langle\mathsf{M}(f)h,g\rangle$
$\text{and}\quad\langle\mathsf{M}(f)g,g\rangle$ $=\frac{1}{2}\int_{\mathbb{R}^{12}}\kappa ff_{*}|\tilde{\nabla}g|^{2}{\,\rm d}\eta\geq 0$
@@
    $\mathsf{M}(f){\,\rm d}\mathsf{S}(f)$ $=\frac{1}{2}\tilde{\nabla}\cdot\Big{(}\kappa ff_{*}\tilde{\nabla}\log f\Big{)}.$
It remains to show the non-interaction conditions. We have, by the anti-symmetry of $\mathsf{J}$ and the definition of $\tilde{\nabla}$:
@@
$\text{and}\quad\mathsf{M}(f){\,\rm d}\mathsf{E}(f)$ $=-\frac{1}{2}\tilde{\nabla}\cdot\Big{(}\kappa ff_{*}\tilde{\nabla}\frac{|v|^{2}}{2}\Big{)}=0.$
As mentioned in the introduction, by taking $x=x_{*}$, we formally obtain the GENERIC formulation for the classical inhomogeneous (kinetic) Landau equation. It is worth mentioning that the classical kinetic Boltzmann equation can also be formally cast into the GENERIC framework  [ottinger1997generic, grmela2018generic].
[…]
H. Duong, School of Mathematics, University of Birmingham, UK
E-mail: h.duong@bham.ac.uk
Z. He, Fakultät für Mathematik, Universität Bielefeld, Postfach 100131, 33501 Bielefeld, Germany
E-mail: zihui.he@uni-bielefeld.de

### v2

@@ 3.3 Variational characterisation for the fuzzy Landau equation
$\int_{\mathbb{R}^{2d}}|{\min\{f\log f,0\}}|{\,\rm d}x{\,\rm d}v$
$\leq{}$ $C\int_{\mathbb{R}^{2d}}f^{1-\kappa}{\,\rm d}x{\,\rm d}v$
$\leq{}$ $C\Big(\int_{\mathbb{R}^{2d}}\big(\langle x\rangle^{2}+\langle v\rangle^{2}\big)^{-\frac{1-\kappa}{\kappa}}{\,\rm d}x{\,\rm d}v\Big)^{\kappa}\Big(\int_{\mathbb{R}^{2d}}\big(\langle x\rangle^{2}+\langle v\rangle^{2}\big)f{\,\rm d}x{\,\rm d}v\Big)^{1-\kappa}$
where we choose $\kappa>0$ small such that $(\langle x\rangle^{2}+\langle v\rangle^{2})^{-\frac{1-\kappa}{\kappa}}\in L^{1}(\mathbb{R}^{2d})$. Hence, we have
@@
$\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\log(f_{t})U_{t}{\,\rm d}\eta{\,\rm d}t\geq-\frac{1}{2}\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t-\frac{1}{2}\int_{0}^{T}\mathcal{A}(f_{t},U_{t}){\,\rm d}t.$
@@
$\mathcal{H}(f_{T})-\mathcal{H}(f_{0})=\frac{1}{2}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\log(f_{t})U_{t}{\,\rm d}\eta{\,\rm d}t$
$\geq-\frac{1}{4}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\big(-\widetilde{\nabla}\log(f_{t})\big)^{2}ff_{*}\kappa{\,\rm d}\eta{\,\rm d}t-\frac{1}{4}\int_{0}^{T}\int_{\mathbb{R}^{4d}}\frac{|U_{t}|^{2}}{{ff_{*}\kappa}}{\,\rm d}\eta{\,\rm d}t.$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{4d}}\big|\widetilde{\nabla}\log(f_{t})\sqrt{ff_{*}\kappa}+\frac{U_{t}}{\sqrt{ff_{*}\kappa}}\big|^{2}{\,\rm d}\eta{\,\rm d}t=0.$
@@
$U=-ff_{*}\kappa\widetilde{\nabla}\log f$
almost everywhere on $0,T]\times\mathbb{R}^{4d}$. The weak formulation (2.13) for the fuzzy Landau equation holds since $(f,-\kappa ff_{*}\widetilde{\nabla}\log f)\in\mathcal{TGRE}_{T}$.
If $f_{t}$ is a $\mathcal{H}$-solutions of (1.1), (notice that the entropy dissipation $\int_{0}^{T}\mathcal{D}(f_{t}){\,\rm d}t<+\infty$ by definition), we take $U=-\kappa ff_{*}\widetilde{\nabla}\log f$, then $(f_{t},U_{t})\in\mathcal{TGRE}_{T}$ and
@@
4 Further discussion and outlook
In this section, we provide further discussion on the GENERIC structure of the fuzzy Landau equation and on possible directions for future works.
4.1 Other GENERIC building blocks
It has been known that an evolution equation may possess different GENERIC structures [Mie23]. For instance, the equation $\partial_{t}\rho=\Delta\rho$ can be written as a gradient flow (that is a GENERIC system in the absence of the reversible component) in at least two different ways
$\partial_{t}\rho=\Delta\rho=\mathsf{M}_{1}(\rho){\,\rm d}\mathsf{S}_{1}(\rho)=\mathsf{M}_{2}(\rho){\,\rm d}\mathsf{S}_{2}(\rho),$
where
$\mathsf{M}_{1}(\rho)\xi=-\mathrm{div}(\rho\nabla\xi),~\mathsf{S}_{1}(\rho)=\int\rho\log\rho$
$\text{and}\quad\mathsf{M}_{2}(\rho)\xi=-\mathrm{div}(\rho^{2}\nabla\xi),~\mathsf{S}_{2}(\rho)=\int\log\rho.$
These two gradient flow structures are respectively related to (via large deviation principle) microscopic models of diffusion and heat conduction processes [MPR14]. The Boltzmann equation can also be formulated in two different GENERIC structures as in [Grm18] and [Ött97].
We now formally show that the (fuzzy) Landau equation also have different GENERIC structures. In fact, the definition of the fuzzy Landau gradient (1.11)
$\widetilde{\nabla}f=\sqrt{A(v-v_{*})}\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}f-\nabla_{v_{*}}f_{*}\big)$
can be generalised to the following families of gradient
$\widetilde{\nabla}_{\xi}f=\xi(|v-v_{*}|)\Pi_{(v-v_{*})^{\perp}}\big(\nabla_{v}f-\nabla_{v_{*}}f_{*}\big)$
$\text{and}\quad\widetilde{\nabla}_{\Xi}f=\Xi(v-v_{*})\big(\nabla_{v}f-\nabla_{v_{*}}f_{*}\big),$
where
$\xi=\xi(|v-v_{*}|):\mathbb{R}_{+}\to\mathbb{R}_{+}\quad\text{and}$
$\Xi=\Xi(v-v_{*}):\mathbb{R}^{d}\to\mathbb{R}^{d\times d}\quad\text{such that}\quad\Xi=\Xi^{T}.$
Moreover, there exists $\xi_{i}=\xi_{i}(v-v_{*}):\mathbb{R}^{d}\to\mathbb{R}_{+}$, $i=1,2$, such that
$\xi^{2}(|v-v_{*}|)\xi_{1}(v-v_{*})=A(v-v_{*}),$
$\xi_{2}(v-v_{*})\Xi^{T}\Pi_{(v-v_{*})^{\perp}}\Xi=A(v-v_{*})\Pi_{(v-v_{*})^{\perp}}.$
Then the corresponding divergence operators, $\widetilde{\nabla}_{\xi}\cdot$ and $\widetilde{\nabla}_{\Xi}\cdot$, are given by
$\widetilde{\nabla}_{\xi}\cdot G=\nabla_{v}\cdot\int_{\mathbb{R}^{2d}}\xi_{1}(|v-v_{*}|)\Pi_{(v-v_{*})^{\perp}}(G-G_{*}){\,\rm d}x_{*}{\,\rm d}v_{*},$
$\widetilde{\nabla}_{\Xi}\cdot G=\nabla_{v}\cdot\int_{\mathbb{R}^{2d}}\Xi(v-v_{*})^{T}(G-G_{*}){\,\rm d}x_{*}{\,\rm d}v_{*},$
for $G:\mathbb{R}^{4d}\to\mathbb{R}^{d}$. Notice that one can also include the spatial kernel $\kappa(x-x_{*})$ into the above gradient operators. For simplicity of notations, we take $\kappa=1$ in the following of this section.
Let $\widetilde{\nabla}_{\xi/\Xi}$ denote the Landau gradient operator $\widetilde{\nabla}_{\xi}$ or $\widetilde{\nabla}_{\Xi}$. The fuzzy Landau equation (1.1) admits infinite many GENERIC building blocks $\{\mathsf{M}_{\xi/\Xi},\mathsf{L},\mathsf{E},\mathsf{S}\}$ where $\mathsf{S}$, $\mathsf{E}$ and $\mathsf{L}$ are given as in Section 1.3, and the operator $\mathsf{M}_{\xi/\Xi}$ is replaced by
$\mathsf{M}_{\xi}(f)g=-\frac{1}{2}\widetilde{\nabla}_{\xi}\cdot\big(\xi_{1}ff_{*}\widetilde{\nabla}_{\xi}g\big)\quad\text{and}$
$\mathsf{M}_{\Xi}(f)g=-\frac{1}{2}\widetilde{\nabla}_{\Xi}\cdot\big(\xi_{2}ff_{*}\Pi_{(v-v_{*})^{\perp}}\widetilde{\nabla}_{\Xi}g\big)$
for all $g\in\mathsf{Z}$. The building block associated to $\mathsf{M}_{\Xi}$ has been proposed in [GGPTZ25].
The variational characterisation in Theorem 1.1 holds for building blocks $\{\mathsf{M}_{\xi},\mathsf{L},\mathsf{E},\mathsf{S}\}$ and $\{\mathsf{M}_{\Xi},\mathsf{L},\mathsf{E},\mathsf{S}\}$ at least on a formal level. To rigorously establish the GENERIC structure, the key step is to prove the chain rule
$\frac{d}{dt}\mathcal{H}(\mu_{t})=\frac{1}{2}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}_{\xi/\Xi}\log f_{t}{\,\rm d}\mathcal{U}_{t}.$
In this article, we have shown the chain rule for $\widetilde{\nabla}_{\xi}$ for the case of $\xi=\sqrt{A}$. Other possible approach is to follow the strategy developed for the Boltzmann equation [Erb23, EH25]. This relies on the commutation property $\widetilde{\nabla}_{\xi},M_{\beta}*_{v}]=0$, that is,
$\big(\widetilde{\nabla}_{\xi}\cdot U\big)*_{v}M_{\beta}=\widetilde{\nabla}_{\xi}\cdot\big(U*_{v,v_{*}}M_{\beta}\big).$
However, it is unclear whether such identity holds in general. This identity may hold for $\widetilde{\nabla}_{\Xi}$ in special cases; for instance, one can take $\Xi=\operatorname{Id}$. Nevertheless, in order to control the collision term, we lack of estimates in the direction of $v-v_{*}$.
4.2 Other kinetic models
@@
Together with previous works [CDDW24, Erb23, EH25] on the Boltzmann/Landau equations, our results cast the spatial homogeneous, fuzzy (delocalised), and classical (at least formally) versions of these equations into a common variational framework. This paves the way for deriving one equation from another by passing to the limit a certain parameter (such as from the Boltzmann equations to the Landau equations in the grazing limit, or from the fuzzy models to the classical ones in the regularisation limits) by generalising the well-established evolutionary $\Gamma$-convergence methods for gradient flows [SS04, Mie16] to GENERIC systems. Another interesting direction for future work is to generalise the results of this paper to other kinetic models such as the relativistic/quantum Boltzmann/Landau equations, and their multi-species models.
@@ Appendix A GENERIC formulation
We consider the space $\mathsf{Z}$ to be the space that consists of all Schwartz functions $f\in\mathcal{S}(\mathbb{R}^{2d})$ such that $f\mathcal{L}\in\mathcal{P}(\mathbb{R}^{2d})$. The space is endowed with the $L^{2}$-inner product $\langle f,g\rangle=\int_{\mathbb{R}^{2d}}fg{\,\rm d}x{\,\rm d}v$. We formally consider the tangent and co-tangent space at $f\in\mathsf{Z}$ as $T_{f}\mathsf{Z}=\{\xi\in L^{2}(\mathbb{R}^{2d})\mid\int_{\mathbb{R}^{2d}}\xi{\,\rm d}x{\,\rm d}v=0\}$ and $T_{f}^{*}\mathsf{Z}=L^{2}(\mathbb{R}^{2d})\}$.
@@
$\mathsf{E}(f)=\int_{\mathbb{R}^{2d}}\frac{|v|^{2}}{2}f{\,\rm d}x{\,\rm d}v\quad\text{and}\quad\mathsf{S}(f)=-\mathcal{H}(f).$
@@
$\mathsf{M}(f)g$ $=-\frac{1}{2}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}g\Big)\quad\text{and}$
$\mathsf{L}(f)g$ $=-\nabla\cdot(f\mathsf{J}\nabla g),\quad\mathsf{J}=\begin{pmatrix}0&\mathsf{id}_{d}\\ -\mathsf{id}_{d}&0\end{pmatrix}$
@@
$\langle\mathsf{L}(f)g,g\rangle=\int_{\mathbb{R}^{2d}}f\nabla g\mathsf{J}\nabla g=0$
@@
$\langle\mathsf{M}(f)g,h\rangle$ $=-\frac{1}{2}\int_{\mathbb{R}^{4d}}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}g\Big)h{\,\rm d}\eta$
$=\frac{1}{2}\int_{\mathbb{R}^{4d}}\kappa ff_{*}\widetilde{\nabla}g\cdot\widetilde{\nabla}h{\,\rm d}\eta=\langle\mathsf{M}(f)h,g\rangle$
$\text{and}\quad\langle\mathsf{M}(f)g,g\rangle$ $=\frac{1}{2}\int_{\mathbb{R}^{4d}}\kappa ff_{*}|\widetilde{\nabla}g|^{2}{\,\rm d}\eta\geq 0$
@@
    $\mathsf{M}(f){\,\rm d}\mathsf{S}(f)$ $=\frac{1}{2}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}\log f\Big).$
It remains to show the non-interaction conditions. We have, by the anti-symmetry of $\mathsf{J}$ and the definition of $\widetilde{\nabla}$:
@@
$\text{and}\quad\mathsf{M}(f){\,\rm d}\mathsf{E}(f)$ $=-\frac{1}{2}\widetilde{\nabla}\cdot\Big(\kappa ff_{*}\widetilde{\nabla}\frac{|v|^{2}}{2}\Big)=0.$
As mentioned in the introduction, by taking $x=x_{*}$, we formally obtain the GENERIC formulation for the classical inhomogeneous (kinetic) Landau equation. It is worth mentioning that the classical kinetic Boltzmann equation can also be formally cast into the GENERIC framework  [Ött97, Grm18].